The inner tube of a Pyrex glass mercury thermometer has a diameter of 0.120 mm. The bulb at the bottom of the thermometer contains 0.420 cm3 of mercury. How far will the thread of mercury move for an increase of 1.00°C? Remember to take into account the expansion of the glass. Coefficient of Linear Expansion of Glass(Pyrex) is 3.25 × 10−6 K−1 and Coefficient of Volume Expansion of Mercury is 182 × 10−6 K−1.


1
Expert's answer
2022-05-22T14:39:07-0400

"h=\frac{4V(1+\beta t)}{\pi d^2(1+\alpha t)^2}=0.9~cm."